# PLAME: Lightweight MSA Design Advances Protein Folding From Evolutionary Embeddings

### v2

@@ Appendix B Training and Sampling Details
Appendix C Extensive Benchmark Results
C.1 Comparison to AI-based MSA Retrieval Methods
We further compare the folding enhancement by MSAs retrieved by AI-based approaches [19]. Without filtering, DHR’s performance is generally higher than AF2 MSA, EvoDiff, and MSAGPT, demonstrating the effectiveness of applying evolutionary embeddings like ESM2.
Zero/Few pLDDT GDT TMscore RMSD LDDT
DHR Zero 63.64 0.51 0.52 12.04 0.55
DHR Few 62.60 0.52 0.57 11.92 0.59
PLAME Zero 66.54 0.53 0.53 11.48 0.57
PLAME Few 66.08 0.54 0.58 12.14 0.60
Table 3: Comparison of folding enhancement to AI-based MSA retrieval method.
C.2 Comparison on Inference Speed and Memory Usage
@@
C.3 Ablation on HiFiAD Filtering
To demonstrate the effectiveness of HiFiAD, we conducted ablation experiments on filtering methods for other baselines. Among the benchmarks, PLAME maintains SOTA performance, and DHR-HiFiAD and PLAME show comparable performance across multiple metrics. Several metrics of EvoDiff, MSAGPT, and DHR (AI-based MSA retrieval methods) all demonstrate improvement after HiFiAD filtering compared to their original filtering methods, which directly proves the benefit of HiFiAD rules for MSA in AF-series structure prediction. This insight can help future work develop higher-quality MSA sequences.
Zero/Few pLDDT GDT TMscore RMSD LDDT
DHR-HiFiAD Zero 66.01 0.53 0.53 11.48 0.57
DHR-HiFiAD Few 66.08 0.55 0.60 12.14 0.60
EvoDiff-HifiAD Zero 58.24 0.46 0.46 13.74 0.51
EvoDiff-HifiAD Few 60.89 0.49 0.54 12.39 0.56
MSAGPT-HifiAD Zero 60.16 0.48 0.48 12.54 0.53
MSAGPT-HifiAD Few 62.63 0.52 0.57 12.18 0.59
Table 5: Ablation on HiFiAD filtering.
C.4 Ablation on Protein Length
We listed the performance of PLAME and AF2 MSA in different length ranges, and found that PLAME shows overall improvement, with the largest improvement in the 100-300 range, followed by the >>300 and <<100 ranges. We believe this is due to our MSA training data being mainly concentrated in the 100-300 length range.
Length Range pLDDT GDT TMscore RMSD LDDT
AF2 MSA <<100 71.03 0.64 0.52 7.77 0.61
AF2 MSA 100-300 59.50 0.49 0.53 12.46 0.54
AF2 MSA >>300 56.29 0.43 0.51 15.67 0.53
PLAME <<100 74.12 0.63 0.52 7.49 0.61
PLAME 100-300 65.55 0.53 0.58 11.58 0.58
PLAME >>300 58.31 0.45 0.53 16.16 0.54
Table 6: Ablation on protein length.
Appendix D Extensive Case Studies
D.1 Case Study on Successful Designs
To further explore the key pattern of the MSA augmentation, we provide a series of sequence and structure visualization in Appendix G. We select representative cases collected from different datasets and range from different lengths to comprehensive evaluate the samples.
@@ D.1 Case Study on Successful Designs
D.2 AlphaFold3 Folding Experiment
To further validate PLAME is an universal MSA generation framework for different folding software, we conduct extensive experiment on AlphaFold3 [5]. We use the same generated MSAs filtered by HiFiAD, augmenting the original MSAs for folding enhancement.
pLDDT GDT TMscore RMSD LDDT pTM
AF2 MSA 68.872 0.578 0.596 10.818 0.617 0.529
PLAME 70.887 0.578 0.595 10.740 0.623 0.539
Table 7: Comparison of folding enhancement based on AlphaFold3
Results follow the same trend as AF2 Mode 3. The stronger the folding baseline, the smaller the performance gain. Improvements are mainly in pLDDT and RMSD, indicating our enhanced MSA primarily improves local structural regions while preserving the global protein architecture.
D.3 Folding Enhancement on Average Proteins
@@
D.4 Further Ablation on MSA Filtering
@@
D.5 Failure Case Analysis
@@
Among the visualized MSA cases, we observed that generated MSAs contained extremely similar sequences (>90% similarity). Specifically, these high-similarity sequences caused all sites to appear more conserved, resulting in a lack of covariation patterns necessary for AlphaFold2 to infer structural contacts. This pattern was evident across all four cases. Notably, for 3bog_B and 8gzu_T3, the generated high-similarity MSAs further enhanced the conservation of already conserved regions, which consequently led to improvements in global metrics.
D.6 De novo Protein Folding Enhancement
We conduct further experiments on De Novo protein cases, where almost of them are orphan. Examples of de novo proteins include 8SK7 (RFDiffusion), 8TNM/8TNO (Chroma), and 8CYK (ProteinMPNN). We followed the same augmentation pattern as the main experiment.
@@ D.6 De novo Protein Folding Enhancement
From Table 11, we observed that PLAME experiences a slight decrease in pLDDT scores while simultaneously showing improvements in other metrics. The generated MSA visualizations in Figures 4 and 5 reveal that most generated sequences maintain $70\%$ similarity to the query sequences. This phenomenon may be attributed to these test cases being highly Out-Of-Distribution (OOD) relative to the training dataset. Nevertheless, the diverse sampling strategy still effectively enhances the profile information of orphan proteins, resulting in substantial performance improvements. Furthermore, we visualized specific local regions where PLAME achieves superior alignment performance as measured by TMscore. Analysis revealed that across all augmented profiles, these high-performing local regions exhibit remarkable conservation, suggesting a strong correlation between sequence conservation patterns and structural alignment quality.
@@
Appendix E Discussion
E.1 Limitations
@@
E.2 Social Impact
PLAME advances multiple sequence alignment (MSA) generation for proteins, offering significant potential benefits for protein structure prediction and engineering applications. This capability could accelerate therapeutic protein development, enhance drug design processes, and facilitate the exploration of novel functional protein sequences. However, we acknowledge important limitations and risks associated with this technology. The model may occasionally generate hallucinated sequences that could mislead downstream protein design efforts if not properly validated. Furthermore, like many powerful biotechnology tools, there exists potential for misuse in designing harmful biological entities.
Appendix F Structure Comparison Visualization
@@
Appendix G Augmented MSA Visualization
@@
Appendix H Failure Case MSA Visualization

### v3

@@ Appendix B Training and Sampling Details
Appendix C Related Works
Protein Structure Prediction
Protein structure prediction methods fall into three main categories: physics-based, homology-based, and deep learning approaches. Physics-based methods, such as AMBER and CHARMM, use molecular physics and energy optimization to simulate protein folding [10, 20]. While offering detailed folding insights, they are computationally expensive and sensitive to initial conditions, often yielding suboptimal results [21, 22, 23]. Homology modeling tools, like Rosetta and HHpred, use MSAs and evolutionary data to predict structures by refining templates from known experimental structures [9, 11]. These methods perform well with suitable templates but struggle with orphan proteins and low-homology families [8, 1]. Deep learning-based methods, such as AlphaFold2 and OmegaFold, use advanced neural architectures and protein templates to achieve near-experimental accuracy with greater speed and scalability [3, 5, 24]. Despite their success, they still depend on high-quality MSAs and struggle with low-homology proteins.
AlphaFold-based Enhancement
Building on AlphaFold’s success, researchers have developed methods to refine specific modules, aiming to improve accuracy or efficiency. These advancements can be grouped into three main categories. The first category focuses on homology expansion techniques, such as MMSeq2 and DeepMSA2, which expand the evolutionary search space to enhance prediction accuracy. However, these methods often slow down inference despite their modest performance gains [25, 26, 27, 28]. The second category targets search acceleration, with methods like ColabFold and ESMFold bypassing the MSA search process to enhance computational efficiency. However, this speedup often results in incomplete evolutionary data, potentially reducing prediction accuracy [6, 29]. The third category leverages generative models to capture protein homology and augment input data, especially for orphan proteins and low-homology families. While promising in specific scenarios, these models struggle with extremely limited evolutionary signals, and their artificial sequences often deviate from traditional MSA distributions, limiting broader applicability [16, 14, 13, 15].
Appendix D Comparison on Inference Speed and Memory Usage
@@
Appendix E Extensive Case Studies
E.1 Case Study on Successful Designs
To further explore the key pattern of the MSA augmentation, we provide a series of sequence and structure visualization in Appendix H. We select representative cases collected from different datasets and range from different lengths to comprehensive evaluate the samples.
@@ E.1 Case Study on Successful Designs
E.2 Folding Enhancement on Average Proteins
@@
E.3 Further Ablation on MSA Filtering
@@
E.4 Failure Case Analysis
@@
Among the visualized MSA cases, we observed that generated MSAs contained extremely similar sequences (>>90% similarity). Specifically, these high-similarity sequences caused all sites to appear more conserved, resulting in a lack of covariation patterns necessary for AlphaFold2 to infer structural contacts. This pattern was evident across all four cases. Notably, for 3bog_B and 8gzu_T3, the generated high-similarity MSAs further enhanced the conservation of already conserved regions, which consequently led to improvements in global metrics.
E.5 De novo Protein Folding Enhancement
We conduct further experiments on De Novo protein cases, where almost of them are orphan. Examples of de novo proteins include 8SK7 (RFDiffusion [38]), 8TNM/8TNO (Chroma [39]), and 8CYK (ProteinMPNN [40]). We followed the same augmentation pattern as the main experiment.
@@ E.5 De novo Protein Folding Enhancement
From Table 8, we observed that PLAME experiences a slight decrease in pLDDT scores while simultaneously showing improvements in other metrics. The generated MSA visualizations in Figures 5 and 6 reveal that most generated sequences maintain $70\%$ similarity to the query sequences. This phenomenon may be attributed to these test cases being highly Out-Of-Distribution (OOD) relative to the training dataset. Nevertheless, the diverse sampling strategy still effectively enhances the profile information of orphan proteins, resulting in substantial performance improvements. Furthermore, we visualized specific local regions where PLAME achieves superior alignment performance as measured by TMscore. Analysis revealed that across all augmented profiles, these high-performing local regions exhibit remarkable conservation, suggesting a strong correlation between sequence conservation patterns and structural alignment quality.
@@
Appendix F Discussion
F.1 Limitations
@@
Appendix G Structure Comparison Visualization
@@
Appendix H Augmented MSA Visualization
@@
Appendix I Failure Case MSA Visualization